# GraphCogent: Overcoming LLMs’ Working Memory Constraints via Multi-Agent Collaboration in Complex Graph Understanding

### v1

@@ 3.1 Sensory Module
$\text{GEC}=\text{Edit\_distance}(G,G^{\prime})\times\frac{\sum_{i=1}^{N}(T_{\text{input}}^{i}+T_{\text{output}}^{i})}{T_{\text{max}}},$ (1)
where $\text{Edit\_distance}(G,G^{\prime})$ measures the structural deviation between the original graph $G$ and the sampled graph $G^{\prime}$; $\sum_{i=1}^{N}(T_{\text{input}}^{i}+T_{\text{output}}^{i})$ quantifies the total token cost of LLM interactions, normalized by $T_{\text{max}}$ (maximum allowed tokens per LLM interaction). The algorithm for calculating Edit_distance and the rationale for selecting token count as the Cost metric are detailed in Appendix E.
@@
Guided by GEC, the Sensory Agent decomposes large graphs into optimally-sized subgraphs to achieve accurate graph transformation. We employ heuristic prompts (prompt details in Appendix I) to guide the agent outputting standardized adjacency lists regardless of input format (symbolic notations, linguistic descriptions, etc.). To handle both weighted and unweighted graph adjacency lists, we implement dual regular expression templates to automatically match these two data formats. These templates precisely extract edge information from the parsed adjacency lists, with the extracted data subsequently validated by the Graph Verifier to ensure structural and semantic accuracy.
Graph Verifier: While GEC optimizes sampling granularity, LLM output uncertainty (shown by non-zero Edit distance in Fig. 4) may cause parsing errors during graph transformation. Since using standard graph data for correction risks information leakage, verification is constrained to known sampling granularity and target formats. The Verifier checks quantity consistency between original and transformed edges, and format compliance with adjacency list standards through regex parsing. Failed transformations trigger Sensory Agent retries.
3.2 Buffer Module
@@
Based on the human episodic buffer mechanism Gong et al. (2024), we construct the Buffer Module to serve as an intermediary that not only stores and integrates graph data but also establishes indices for different data formats. This module transforms raw graph text representations into multiple standardized formats: NetworkX for graph algorithms, NumPy for numerical operations, and PyG for tensor requirement tasks.
@@ 3.3 Execution Module
Execution Module tackles code construction complexity and inefficiency challenges through two components: Reasoning Agent and Model Agent. The Reasoning Agent performs task analysis and tool calling using pre-built common toolset. It first evaluates whether a task can be solved using existing tools, then directly executes the appropriate tool when the task falls within the toolset’s coverage. For out-toolset tasks, it activates the Model Agent to generate specialized models. These models operate directly on the Buffer Module’s preprocessed data structures, avoiding the reliability issues of full code generation while maintaining the system’s adaptability to complex requirements.
3.3.1 Reasoning Agent
The Reasoning Agent’s tool calling and task discrimination capabilities are strengthened through a two-stage instruction tuning combining Thinking-Enhanced Supervised Fine-Tuning Hu et al. (2022) (SFT) with Direct Preference Optimization Rafailov et al. (2023) (DPO). This two-stage approach achieves both precise tool selection for in-toolset tasks and reliable discrimination of out-toolset scenarios.
Phase 1: Thinking-Enhanced Supervised Fine-Tuning for Tool Selection. We pair graph reasoning problems Prob with toolset descriptions Toolset, forming composite queries Ques = <Prob, Toolset>. These queries are processed by DeepSeek-R1 to generate outputs Ans = <Think, Tool>, where Think captures the reasoning process to improve decision transparency and tool selection robustness, while Tool indicates the selected tool. We curate high-quality training data by retaining only instances where the tool selection is correct, ensuring SFT process optimizes the following objective:
@@ 3.3.1 Reasoning Agent
$\mathcal{L}_{\text{SFT}}=-\sum_{i=1}^{N}\log P(\langle\textit{Think, Tool}\rangle_{i}|\langle\textit{Prob, Toolset}\rangle_{i};\theta),$ (2)
where $N$ is the number of training examples, and $\theta$ represents the LLMs’ parameters.This phase enables the agent to develop active thinking and tool selection capabilities.
Figure 5: An overview of Reasoning Agent and Model Agent training process.
While Thinking-Enhanced SFT provides basic tool selection capabilities, Appendix C.3 reveals a critical limitation: the agent tends to force-fit out-toolset tasks into existing tools despite its active thinking abilities. This observation motivates our second-phase alignment approach.
Phase 2: Direct Preference Optimization Driven Coverage Recognition. To refine the agent’s tool coverage discrimination ability, we employ DPO to align its reasoning preferences. For each problem instance Ques, we generate multiple candidate responses from the SFT-tuned agent. These responses are then classified into two categories: Preferred responses ($R_{w}$): Correctly identify the out-toolset tasks with valid reasoning; Less preferred responses ($R_{l}$): Incorrectly associate tasks with tools. The DPO objective sharpens the agent’s decision boundary by optimizing:
$\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{(\text{Ques},R_{w},R_{l})}\left[\log\sigma\left(\beta\left(\log\frac{P_{\theta}(R_{w}|\text{Ques})}{P_{\text{SFT}}(R_{w}|\text{Ques})}-\log\frac{P_{\theta}(R_{l}|\text{Ques})}{P_{\text{SFT}}(R_{l}|\text{Ques})}\right)\right)\right],$ (3)
where $\beta$ controls deviation from the SFT-tuned model. This penalizes overconfidence in-toolset coverage and rewards accurate out-toolset detection. For each task, the Reasoning Agent first determines if it can be solved using the available toolset. If solvable, it constructs the tool calling using relevant Buffer Module data formats and executes the computation. Otherwise, it invokes the Model Agent for out-toolset processing.
3.3.2 Model Agent
The Model Agent handles out-toolset tasks by generating task-specific models. This maintains framework’s reliability through direct operation on the Buffer Module’s preprocessed data structures, avoiding full code generation issues. To develop the model generation capability, we construct a dataset consists of classical graph algorithms (e.g., PageRank) and neural models (e.g., GCN, GAT) from GitHub. Using LLMs (DeepSeek-v3 DeepSeek-AI (2025), GPT-4o OpenAI (2023), Claude-3.7 Anthropic (2024)), we generate executable graph reasoning codes. For each successful execution, we extract the core reasoning model component, creating training pairs (Ques, Model) to fine-tune Llama for model generation.
$\mathcal{L}_{\text{SFT}}=-\sum_{i=1}^{N}\log P(\textit{Model}_{i}|\textit{Ques}_{i};\theta),$ (4)
where $N$ represents validated examples and $\theta$ denotes LLMs’ parameters. This phase enables agent to generate task-specific models based on tasks while ensuring compatibility with preprocessed data. For out-toolset tasks from Reasoning Agent, Model Agent generates task-specific models based-on task descriptions and Buffer Module data formats. The framework integrates preprocessed data with generated models into executable code and produces final results.
3.3.3 Training Settings
@@
We use Llama3.1-8B as the backbone for all agents, fine-tuned with LoRA Hu et al. (2022). Both Reasoning Agent and Model Agent are trained on Graph4real dataset. The training set employs 4 of 20 prompt templates, with their corresponding tasks excluded from the test set to ensure fair evaluation.
4 Experiments
4.1 Experimental Settings
@@
Dataset. We evaluate on Graph4real, a real-world benchmark with graphs from 40 to 1,000 nodes spanning transportation, social, web, and citation domains. The dataset includes both in-toolset and out-toolset tasks, with 500 test instances per task and scale (across domains). To ensure fair comparisons, we use standard adjacency lists for graph text representation for main experiments in Table 1. Additional experiments on public datasets are provided in Appendix D.
@@
Baseline Methods. We compare three types of approaches: (1) Text-based methods including GPT-4o OpenAI (2023), Claude-3.7 Anthropic (2024), and DeepSeek-R1 DeepSeek-AI (2025) with 2-shot CoT prompting. (2) Tool-based methods including Graph-Toolformer Zhang (2023), GraphTool-Instruction Wang et al. (2025b) on Llama3.1-8B Touvron et al. (2023) and GPT-4o with Function Calling OpenAI (2023); (3) Agent-based methods featuring the state-of-the-art GraphTeam Li et al. (2024) framework with GPT-4o-mini OpenAI (2023). All LLMs’ versions and settings are presented in Table 6.
@@
Evaluation Metrics. We use accuracy for all tasks except traffic flow prediction (measured by MAE). Detailed metric computations are in Appendix D.
@@
Table 1: Performances of GraphCogent and other baselines on Graph4real in-toolset tasks.
@@
Text GPT-4o OpenAI (2023) 100 22.2 20.6 25.8 77.2 78.2 100 84.2 29.8 59.8
Claude-3.7 Anthropic (2024) 99.2 22.4 21.0 21.2 79.8 81.8 100 80.2 32.2 59.8
DeepSeek-R1 DeepSeek-AI (2025) 100 32.6 26.6 30.8 86.6 86.8 100 90.4 31.0 65.0
Tool Graph-TF Zhang (2023) 48.0 75.6 71.0 63.8 44.2 66.8 50.2 77.2 50.0 60.8
GraphTool-Ins Wang et al. (2025b) 94.2 98.0 97.8 98.2 96.8 99.0 100 100 97.8 98.0
GPT4o-FC OpenAI (2023) 95.0 99.2 98.4 99.0 99.2 100 100 100 99.0 98.9
Agent GraphTeam Li et al. (2024) 86.6 84.8 82.8 80.2 97.6 80.0 96.8 70.8 97.6 86.4
@@
Text GPT-4o OpenAI (2023) 90.4 6.8 5.4 7.2 54.0 67.6 99.8 58.6 6.8 44.1
Claude-3.7 Anthropic (2024) 93.0 6.2 4.8 6.8 55.8 73.0 99.2 52.2 7.2 44.2
DeepSeek-R1 DeepSeek-AI (2025) 98.0 7.4 6.2 8.0 60.6 76.6 100 60.4 7.4 47.2
Tool Graph-TF Zhang (2023) 43.2 68.0 63.8 57.4 39.8 60.0 45.2 69.4 45.0 54.6
GraphTool-Ins Wang et al. (2025b) 84.8 89.0 88.0 88.4 89.0 89.0 90.6 92.0 88.0 88.8
GPT4o-FC OpenAI (2023) 85.4 89.2 88.6 89.0 89.2 89.4 97.0 94.2 89.0 90.1
Agent GraphTeam Li et al. (2024) 91.6 80.2 73.0 65.8 93.2 69.6 89.6 65.4 95.2 80.4
@@
Note: Graph-TF, GraphTool-Ins ,and GPT4o-FC are abbreviations for Graph-Toolformer, GraphTool-Instruction, and GPT4o-Function Calling.
4.2 Main Result
Figure 6: Code execution rate on out-toolset tasks.
The performance results are reported in Table 1. Our method significantly outperforms all baseline methods across toolset-covered tasks. As for text-based methods, they demonstrate advantages on small-scale simple Structural Querying tasks (e.g., DeepSeek-R1 achieves 100% on edge existence). However, these methods show significant performance degradation when dealing with problems requiring multi-step reasoning (e.g., only 30.8% accuracy on shortest path), and this issue becomes more pronounced as the scale increases. Tool-based approaches such as GPT4o-FC demonstrates robust performance across all tasks on small-scale graphs by leveraging GPT4o’s powerful reasoning capabilities. However, they suffer 10% performance degradation when scaling from 40 to 100 nodes. The agent-based GraphTeam demonstrates limitations in code generation across multiple tasks. The complex data preprocessing and reasoning requirements overload the model’s working memory, resulting in lower performance. In contrast, our method maintains state-of-the-art performance across all scales with less than 1% variance. Notably, while no baseline could handle large-scale graphs, we have only included the results of our approach.
Table 2: Performance evaluation on out-toolset tasks.
Transportation Social Web Citation
Method Max Diameter Traffic Maxcore Connect. Link Common PageRank Link Reference Cluster. Node
Flow Calcu. Pred. Calcu. Compo. Pred. Neighbor Calcu. Pred. Match Coeff. Class.
GraphTeam 72.4 61.6 96.7 69.8 72.0 76.4 51.2 76.8 71.6 78.8 76.4 72.0
GraphCogent 89.6 92.4 35.1 91.6 87.6 85.2 88.8 92.6 86.4 90.0 85.2 89.6
Note: Except for the traffic prediction (which values represent MAE), all other values represent Accuracy.
Table 3: Reasoning token comparison between GraphTeam and GraphCogent.

### v2

@@ 4.2. Sensory Module
(1) $\text{GEC}=\text{Edit\_distance}(G,G^{\prime})\times\frac{\sum_{i=1}^{N}(T_{\text{input}}^{i}+T_{\text{output}}^{i})}{T_{\text{max}}},$
where $\text{Edit\_distance}(G,G^{\prime})$ measures the structural deviation between the original graph $G$ and the sampled graph $G^{\prime}$; $\sum_{i=1}^{N}(T_{\text{input}}^{i}+T_{\text{output}}^{i})$ quantifies the total token cost of LLM interactions, normalized by $T_{\text{max}}$ (maximum allowed tokens per LLM interaction). The algorithm for calculating Edit_distance and the rationale for selecting token count as the Cost metric are detailed in Appendix C.
@@
Figure 5. Overview of Execution Agent and Tool Creator training. In-toolset task (top) initializes a thinking policy from Think and Tool pairs, applies a CMPO method to refine Execution Agent’s tool capability discrimination; Out-toolset task (bottom) uses the fine-tuned Tool Creator to synthesize minimal task-specific tools, and backfills them into the common toolset.
Guided by GEC, the Sensory Agent decomposes large graphs into optimally sized subgraphs to achieve accurate transformation. We employ heuristic prompts to elicit standardized adjacency lists regardless of input format (e.g., symbolic notations, linguistic descriptions). To handle both weighted and unweighted adjacency lists, we implement dual regular-expression templates that automatically match these formats; extracted edges are subsequently validated by the Graph Verifier. For text-represented graphs, we implement a sliding-window sampler whose window size is selected by GEC; by default with Llama3.1-8B we use 50 edges per window. Each window carries a fixed 5% edge overlap to preserve cross-window structures at boundaries. For structured graph files (e.g., edgelist, txt, and json), we bypass partitioning and load them directly into the Buffer Module without windowing.
Graph Verifier. While GEC controls the granularity, LLM outputs can still introduce parsing noise during transformation. The Verifier checks quantity consistency ($|V|$, $|E|$) and adjacency-list format compliance via regex parsing, and failed transformations trigger Sensory Agent retries. After per-window transformation, the Verifier merges subgraphs via a deterministic set-union with de-duplication and then runs a quick quantity-and-format check to ensure global consistency before dispatching any reasoning. By default, text graphs use a 50-edge window with 5% overlap, and file-based graphs are not partitioned. Experiments on how the Graph Verifier enhances transformation reliability and ensures global consistency are provided in Appendix B.1.
4.3. Buffer Module
@@
Based on the human episodic buffer mechanism (Gong et al., 2024), we construct the Buffer Module to serve as an intermediary that not only stores and integrates graph data but also establishes indices for different data formats. This module transforms raw graph text representations into multiple standardized formats: NetworkX for graph algorithms, NumPy for numerical operations, and PyG for tensor requirement tasks.
@@ 4.4. Execution Module
Execution Module tackles code construction complexity and inefficiency through two collaborating components: Execution Agent and Tool Creator. The Execution Agent performs task analysis and tool invocation using a pre-built common toolset. It first assesses whether a problem instance is solvable within the toolset’s capabilities via invoking existing tools; when covered, it directly executes the appropriate tool. For cases beyond current capabilities, it issues a creation request to the Tool Creator, which synthesizes a minimal task-specific tool and fills the pending invocation. By directly operating on the Buffer Module’s preprocessed data results, it avoids the fragility of full-code generation while retaining adaptability to complex requirements. Newly created tools can also be registered back into the toolset for future reuse.
4.4.1. Execution Agent
We adopt a two-stage instruction-tuning pipeline that couples (i) thinking policy initialization for robust tool invocation with (ii) capability-margin preference optimization (CMPO) for reliable tool capability discrimination. The first stage follows instruction-tuned policy learning with explicit thinking traces (Hu et al., 2022); the second stage follows preference-based policy learning (Rafailov et al., 2023) to shape a likelihood-ratio margin around the tool capability boundary.
Phase 1: Thinking Policy Initialization for Tool Invocation.
We pair graph reasoning problems Prob with toolset descriptions Toolset to form composite queries $\textit{Ques}=\langle\textit{Prob},\textit{Toolset}\rangle$. These are processed by DeepSeek-R1 to produce $\textit{Ans}=\langle\textit{Think},\textit{Tool}\rangle$, where Think exposes the reasoning trace to enhance routing stability and auditability, and Tool indicates the selected tool. We curate pairs with correct invocation and optimize the maximum-likelihood objective:
@@ 4.4.1. Execution Agent
(2) $\mathcal{L}_{\text{init}}=-\sum_{i=1}^{N}\log P_{\theta}\!\big(\langle\textit{Think},\textit{Tool}\rangle_{i}\,\big|\,\langle\textit{Prob},\textit{Toolset}\rangle_{i}\big),$
with the factorization:
$\log P_{\theta}(\textit{Ans}\mid\textit{Ques})=\underbrace{\log P_{\theta}(\textit{Think}\mid\textit{Ques})}_{\text{thinking generation}}+\underbrace{\log P_{\theta}(\textit{Tool}\mid\textit{Think},\textit{Ques})}_{\text{tool selection}}.$
While the first stage provides strong invocation priors, Appendix B.3 shows a boundary effect: the agent may still over-rely on existing tools for tasks that actually require creation. This motivates a decision margin for the model’s tool capability. For later unification, we define a policy score relative to a frozen reference policy $P_{\text{ref}}$ for subsequent preference optimization and tool-capability discrimination:
(4) $s_{\theta}(y\mid x)=\log P_{\theta}(y\mid x)-\log P_{\text{ref}}(y\mid x).$
Phase 2: Capability-Margin Preference Optimization.
To refine agent’s tool capability discrimination, we adopt following two strategies.
(a) Action-segment scoring. $\mathcal{M}_{\text{tool}}$ denotes the token indices belonging to the action segment (demarcated by special tool markers in data). We compute a policy score centered at a frozen reference policy $P_{\text{ref}}$ but restricted to the action segment:
$s_{\theta}^{\text{tool}}(y\mid x)=\sum_{t\in\mathcal{M}_{\text{tool}}}\!\log P_{\theta}(y_{t}\mid y_{<t},x)-\sum_{t\in\mathcal{M}_{\text{tool}}}\!\log P_{\text{ref}}(y_{t}\mid y_{<t},x).$
The score aggregates only the decision-critical tokens within the tool segment, isolating action selection from reasoning-style variance. We then quantify the pairwise preference between a preferred response $R_{w}$ and a less-preferred response $R_{l}$ by a likelihood-ratio margin:
(6) $\Delta s^{\text{tool}}=s_{\theta}^{\text{tool}}(R_{w}\!\mid x)-s_{\theta}^{\text{tool}}(R_{l}\!\mid x).$
(b) Capability-aware hard negatives. For each $x$, we sample a candidate set $\mathcal{C}_{K}(x)$ from the initialized policy and select a challenging negative via the reference action-segment likelihood:
(7) $R_{l}\;=\;\operatorname*{argmax}_{y\in\mathcal{C}_{K}(x)\setminus\{R_{w}\}}\ell_{\text{ref}}^{\text{tool}}(y\mid x),$
(8) $\ell_{\text{ref}}^{\text{tool}}(y\mid x)\;=\;\sum_{t\in\mathcal{M}_{\text{tool}}}\log P_{\text{ref}}(y_{t}\mid y_{<t},x).$
This selection aims to identify the most confusing incorrect answer for the agent based on the reference likelihood, forcing the agent to learn to distinguish fine-grained differences between them.
We then optimize a logistic objective with a fixed margin $m$:
(9) $\mathcal{L}_{\text{CMPO}}^{(\text{margin})}=-\mathbb{E}_{(x,R_{w},R_{l})}\Big[\log\sigma\big(\beta\,(\Delta s^{\text{tool}}-m)\big)\Big],$
which encourages the agent not only to prefer the correct tool over the negative, but to maintain a minimum confidence gap between them, thereby improving robustness near the decision boundary.
At execution time, we define a capability score $\gamma(x)=\Delta s^{\text{tool}}$ and employ a two-threshold hysteresis rule to stabilize tool capability discrimination near the boundary, thereby further ensuring the stability and robustness of the agent’s decisions:
(10) $\text{route}(x)=\begin{cases}\mathcal{A}_{\text{invoke}},&\gamma(x)\geq\tau_{\text{high}},\\[2.0pt] \mathcal{A}_{\text{create}},&\gamma(x)\leq\tau_{\text{low}},\\[2.0pt] \text{keep previous decision},&\text{otherwise}.\end{cases}$
4.4.2. Tool Creator
The Tool Creator handles missing-capability cases by generating only the tool body under a prompt template and returning it to fill the pending invocation. The template supplies the task description, admissible libraries and variable names, and the expected return object; outer interface and skeleton code is provided by the framework at runtime. By executing directly on the Buffer Module’s preprocessed data structures, this design keeps generation reliable and avoids brittle full-file synthesis.
We curate a corpus from classical graph algorithms (e.g., PageRank) and neural models (e.g., GCN, GAT) on GitHub. Using LLMs (DeepSeek-v3 (DeepSeek-AI, 2025), GPT-4o (OpenAI, 2023), Claude-3.7 (Anthropic, 2024)), we synthesize code and retain only samples that execute successfully. For each retained sample, we extract the body segment between special body marker. On the input side, we concatenate the original query with its task template to form $\textit{Ques}^{\text{Tpl}}=\langle\textit{Ques},\mathcal{T}\rangle$, and keep a minimal contextual scaffold $c_{i}$ (e.g., necessary imports, a function signature, lightweight placeholders) to induce body-only completion.
Tool Creation Fine-tuning. We optimize a body-only maximum likelihood objective:
(11) $\mathcal{L}_{\text{train}}=-\sum_{i=1}^{N}\ \sum_{t=1}^{M_{\text{body},i}}\log P_{\theta}\!\big(y_{i,t}\,\big|\,y_{i,<t},\ \textit{Ques}^{\text{Tpl}}_{i},\ c_{i}\big),$
where $N$ is the number of validated examples, $M_{\text{body},i}$ denotes the number of tokens in the body segment of example $i$ (we reindex body tokens as $t=1,\dots,M_{\text{body},i}$), $y_{i,t}$ is the $t$-th target token within the body. This objective focuses learning on the implementation logic while leaving interface details to the framework. At inference, the template provides the necessary task and contextual cues; the Tool Creator emits the body segment to satisfy the Execution Agent’s pending invocation, which the framework assembles with the runtime skeleton and executes to produce final results.
4.4.3. Training Settings
@@
We use Llama3.1-8B as the backbone for all agents. For capability-aware negatives, the candidate set size is $K=8$. Two-thresholds are set to $\tau_{\text{low}}=0.55$ and $\tau_{\text{low}}=0.45$. Both Execution Agent and Tool Creator are trained on Graph4real dataset. The training set employs 4 of 20 prompt templates, with their corresponding tasks excluded from the test set to ensure fair evaluation. More detailed settings can be founed in Appendix A.
5. Experiments
5.1. Experimental Settings
@@
Dataset. We evaluate on Graph4real, a real-world benchmark with graphs from 40 to 1,000 nodes spanning transportation, social, web, and citation domains. The dataset includes both in-toolset and out-toolset tasks, with 500 test instances per task and scale (across domains). To ensure fair comparisons, we use standard adjacency lists for graph text representation for main experiments in Table 1. Additionally, we introduce public datasets: NLGraph (Wang et al., 2023), Talk like a Graph (Fatemi et al., 2024), GraphWiz (Chen et al., 2024), GraphArena (Tang et al., 2025), GraCoRe (Yuan et al., 2025) and GraphInstruct (Luo et al., 2024).
@@
Baseline Methods. We compare three types of methods: (1) Text-based methods including GPT-4o (OpenAI, 2023), Claude-3.7 (Anthropic, 2024), and DeepSeek-R1 (DeepSeek-AI, 2025) with 2-shot CoT prompting. (2) Tool-based methods including Graph-Toolformer (Zhang, 2023), GraphTool-Instruction (Wang et al., 2025) on Llama3.1-8B (Touvron et al., 2023) and GPT-4o with Function Calling (OpenAI, 2023); (3) Agent-based methods are the state-of-the-art GraphTeam (Li et al., 2024) framework with GPT-4o-mini (OpenAI, 2023) and reproduced GCoder (Zhang et al., 2024) with Llama3.1-8b (Touvron et al., 2023). All LLMs’ versions and settings are in Appendix A.
@@
Evaluation Metrics. We use accuracy for all tasks except traffic flow prediction (measured by MAE).
@@
Table 1. Performances of GraphCogent and other baselines on Graph4real in-toolset tasks. Note: Graph-TF, GraphTool-Ins ,and GPT4o-FC are abbreviations for Graph-Toolformer, GraphTool-Instruction,and GPT4o-Function Calling.
@@
Text GPT-4o (OpenAI, 2023) 100 22.2 20.6 25.8 77.2 78.2 100 84.2 29.8 59.8
Claude-3.7 (Anthropic, 2024) 99.2 22.4 21.0 21.2 79.8 81.8 100 80.2 32.2 59.8
DeepSeek-R1 (DeepSeek-AI, 2025) 100 32.6 26.6 30.8 86.6 86.8 100 90.4 31.0 65.0
Tool Graph-TF (Zhang, 2023) 48.0 75.6 71.0 63.8 44.2 66.8 50.2 77.2 50.0 60.8
GraphTool-Ins (Wang et al., 2025) 94.2 98.0 97.8 98.2 96.8 99.0 100 100 97.8 98.0
GPT4o-FC (OpenAI, 2023) 95.0 99.2 98.4 99.0 99.2 100 100 100 99.0 98.9
Agent GraphTeam (Li et al., 2024) 86.6 84.8 82.8 80.2 97.6 80.0 96.8 70.8 97.6 86.4
GCoder (Zhang et al., 2024) 82.0 84.6 83.4 82.8 90.6 80.6 88.0 74.0 87.8 83.8
@@
Text GPT-4o (OpenAI, 2023) 90.4 6.8 5.4 7.2 54.0 67.6 99.8 58.6 6.8 44.1
Claude-3.7 (Anthropic, 2024) 93.0 6.2 4.8 6.8 55.8 73.0 99.2 52.2 7.2 44.2
DeepSeek-R1 (DeepSeek-AI, 2025) 98.0 7.4 6.2 8.0 60.6 76.6 100 60.4 7.4 47.2
Tool Graph-TF (Zhang, 2023) 43.2 68.0 63.8 57.4 39.8 60.0 45.2 69.4 45.0 54.6
GraphTool-Ins (Wang et al., 2025) 84.8 89.0 88.0 88.4 89.0 89.0 90.6 92.0 88.0 88.8
GPT4o-FC (OpenAI, 2023) 85.4 89.2 88.6 89.0 89.2 89.4 97.0 94.2 89.0 90.1
Agent GraphTeam (Li et al., 2024) 91.6 80.2 73.0 65.8 93.2 69.6 89.6 65.4 95.2 80.4
GCoder (Zhang et al., 2024) 80.0 68.0 60.8 68.2 82.4 71.4 78.2 67.4 82.8 73.2
@@
Table 2. Performance evaluation on out-toolset tasks. Note: Except for the traffic prediction (which values represent MAE), all other values represent Accuracy.
Web Transportation Social Citation
Method Common PageRank Link Max Diameter Traffic Maxcore Connect. Link Reference Cluster. Node
Neighbor Calcu. Pred. Flow Calcu. Pred. Calcu. Compo. Pred. Match Coeff. Class.
GraphTeam 51.2 76.8 71.6 72.4 61.6 96.7 69.8 72.0 76.4 78.8 76.4 72.0
GCoder 49.0 74.0 68.5 70.0 58.8 99.2 67.2 69.0 72.0 75.0 74.0 68.0
GraphCogent 88.8 92.6 86.4 89.6 92.4 35.1 91.6 87.6 85.2 90.0 85.2 89.6
5.2. Main Result
The performance results are reported in Table 1. Our method significantly outperforms all baseline methods across toolset-covered tasks. As for text-based methods, they demonstrate advantages on small-scale simple Structural Querying tasks (e.g., DeepSeek-R1 achieves 100% on edge existence). However, these methods show significant performance degradation when dealing with problems requiring multi-step reasoning (e.g., only 30.8% accuracy on shortest path), and this issue becomes more pronounced as the scale increases. Tool-based approaches such as GPT4o-FC demonstrates robust performance across all tasks on small-scale graphs by leveraging GPT4o’s powerful reasoning capabilities. However, they suffer 10% performance degradation when scaling from 40 to 100 nodes. The agent-based GraphTeam and GCoder demonstrate limitations in code generation across multiple tasks. The complex data preprocessing and reasoning requirements overload the model’s working memory, resulting in lower performance. In contrast, our method maintains state-of-the-art performance across all scales with less than 1% variance. Notably, while no baseline could handle large-scale graphs, we have only included the results of our approach.
Table 3. Reasoning token comparison between agent baselines. Note: Input and Output are abbreviation for average input tokens and average output tokens respectively.